OpenStudy (anonymous):

Try to solve it.. 5+3= 1628 9+1= 80810 8+6= 28214 5+4= 919 Then 6+2 =....?

OpenStudy (diyadiya):

2847

OpenStudy (anonymous):

6+2 = \(\pi\)

OpenStudy (diyadiya):

pi?

OpenStudy (anonymous):

@Diyadiya: wrong.. its not 2847...

OpenStudy (anonymous):

what technique did you apply?

OpenStudy (diyadiya):

sorry

OpenStudy (anonymous):

How you get 2847 ?

OpenStudy (diyadiya):

its 3248

OpenStudy (vishal_kothari):

3248

OpenStudy (vishal_kothari):

it's right...

OpenStudy (anonymous):

Nice Diya... you are genius ;)

OpenStudy (anonymous):

whats the trick behind it?

OpenStudy (diyadiya):

Not Genius ,as i have seen this question before ,i know the technique :)

OpenStudy (anonymous):

not trick it is a sequence ... isn't it?

OpenStudy (vishal_kothari):

6^2-2^2=32

OpenStudy (vishal_kothari):

6-2=4,6+2=8

OpenStudy (anonymous):

what about 48?

OpenStudy (diyadiya):

last number is obtained by adding the numbers here 6+2=8 Second last number by subtracting the numbers 6-2=4 first two numbers by multiplying 4*8 4*8=32

OpenStudy (anonymous):

Hmmmmmmmmmmmmmmmmmm

OpenStudy (anonymous):

6+2=8 always

OpenStudy (anonymous):

6+2 = 9 idiots.

OpenStudy (anonymous):

can you explain how, mohammed?

OpenStudy (anonymous):

as far as I know, its 8

OpenStudy (anonymous):

that's amazing ! i like it ,3248

OpenStudy (anonymous):

(add*sub)(sub)(add) 5+3=(8*2)(2)(8)=1628 so 6+2=(8*4)(4)(8)=3248

OpenStudy (anonymous):

Answer is 3248... Congratz to them who solve it...:)